# Archetypal SAE: Adaptive and Stable Dictionary Learning for Concept Extraction in Large Vision Models

### v1

@@ E.3 Controlled Rank and Stability in A-SAE Dictionaries
Proposition E.3 (Rank Bound of Archetypal Dictionaries).
@@
E.4 Bounding OOD score with Archetypal Constraints
@@
Proposition E.4 (OOD Measure with Non-Interfering Archetypes).
@@
Appendix F Kernel for JumpReLU
@@
Appendix G Soft Identifiability Benchmark
@@
We generate twelve synthetic datasets using Midjourney API666www.midjourney.com. For each of these datasets, we programmatically create 4,000 images. These images are constructed by collaging four distinct objects selected from the predefined set, such as different types of gems. Each dataset is generated from between 9 and 20 unique objects, with the dictionary size set exactly to the number of true generative factors (unique object number). An example of some datasets used in the benchmark is shown in Figure 14.

### v2

@@ F.3 Controlled Rank and Stability in A-SAE Dictionaries
Proposition F.3 (Rank Bound of Archetypal Dictionaries).
@@
F.4 Bounding OOD score with Archetypal Constraints
@@
Proposition F.4 (OOD Measure with Non-Interfering Archetypes).
@@
Appendix G Kernel for JumpReLU
@@
Appendix H Soft Identifiability Benchmark
@@
We generate twelve synthetic datasets using Midjourney API777www.midjourney.com. For each of these datasets, we programmatically create 4,000 images. These images are constructed by collaging four distinct objects selected from the predefined set, such as different types of gems. Each dataset is generated from between 9 and 20 unique objects, with the dictionary size set exactly to the number of true generative factors (unique object number). An example of some datasets used in the benchmark is shown in Figure 14.